MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2006/11/01
OBSERVER: T.Sumi
CONDITION: slightly cloudy
COMMENT: thin cloud was illuminated by moon, sky was very bright, No good dark was taken after 11 tries.
Compressor gas pressure: MPa
Expected Observation Time: START 21:46 END 05:15


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB4714-472411563360
LMC709-71244613800
SMC250-250192700
ALERT-000
GRB-000
TOTAL---1611119860 (05h31m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
232282329242
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
21122111130


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
333333444433
Field13Field14Field15Field16
3300


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111111000


SPECIAL TARGET NAME

CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
21:40observingstart, delayed due to try to get god dark, but failed
22:00observingslightly cloudy
23:00observingslightly cloudy
00:00observingslightly cloudy
01:00observingslightly cloudy
02:00observingslightly cloudy
03:00observingslightly cloudy
04:00observingslightly cloudy
05:00observingslightly cloudy
05:14observing@stop


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day